OpenStudy (anonymous):

Three consecutive odd integers are such that the square of the second is 96 less than the square of the third. Find the three integers

hero (hero):

(x+2)^2 = (x+4)^2 - 96

OpenStudy (anonymous):

Let the 3 odd integers be n, n+2, n+4 Set up equation (n+2)^2=(n+4)^2-96

hero (hero):

Wow Judith, I'm impressed

OpenStudy (anonymous):

Thank you..

OpenStudy (anonymous):

n^2++4n+4=n^2+8n+16-96 combine like terms 4n=84 n=21 the three integers are 21,23,25

hero (hero):

I agree with Judith

hero (hero):

Judith, did you get any help with the quadratic formula?

OpenStudy (anonymous):

yes i did